# Rethinking Data Augmentation for Robust LiDAR Semantic Segmentation in Adverse Weather

### v1

@@ Methods
Figure 6: Qualitative results of LPD from validation set of SemanticKITTI. It can be observed that the point drop facilitated by the LPD occurs not just locally but extends across all points.
Table 5: Experiments on ASJ/DSJ relative to all-point jittering/CSJ. All models trained on SemanticKITTI and validated on SemanticSTF. The values in parentheses indicate the performance improvement or decrease over the baseline model.
Methods
Clean
D-fog
L-fog
Rain
Snow
mIoU
Baseline 63.9 30.7 30.1 29.7 25.3 31.4
Jittering 61.7 (-2.2) 34.3 36.3 33.2 32.4 36.9 (+5.5)
CSJ 63.2 (-0.7) 32.9 31.1 28.8 23.2 31.2 (-0.2)
ASJ 62.1 (-1.8) 33.3 35.4 37.8 31.6 36.8 (+5.4)
DSJ 62.2 (-1.7) 34.6 36.8 36.1 31.2 38.0 (+6.6)
Analysis of DSJ and ASJ. Our study compared DSJ and ASJ against jittering applied to all points. Additionally, to evaluate Cube-Selective Jittering (CSJ), which involves jittering points within a random cubic area. According to Tab. 5, jittering all points improved robustness in +5.5 mIoU but resulted in a -2.2 mIoU performance decrease on clean data. In contrast, CSJ led to a decrease in mIoU across all adverse weather conditions in a -0.2 mIoU. This outcome suggests that not considering the prior spherical shape of LiDAR data when choosing jittering locations may be inadequate for countering adverse weather. Conversely, ASJ demonstrated improvements (+2.6 mIoU in dense fog, +5.3 mIoU in light fog, +7.9 mIoU in rain, and +6.3 mIoU in snow) with a relatively minor -1.8 mIoU performance decrease in clean data. Also, DSJ showed similar trends with increases in mIoU across various adverse weather conditions and a more moderate -1.7 mIoU drop in clean data performance. These results indicate that ASJ and DSJ effectively simulate coordinate distortion while minimizing the performance gap with clean data.
Table 6: Ratio of points remaining after LPD compared to origianl data. This contrasts with random drop, which exhibits a uniform remaining ratio across different distances.
Distance 0$\sim$10m 10$\sim$20m 20$\sim$30m 30$\sim$40m 40$\sim$50m 50$\sim$60m 60$\sim$70m 70$\sim$80m 80$\sim$90m
Ratio 80.6 78.6 78.7 77.4 76.2 75.5 76.2 76.4 72.5
Analysis of LPD. In our research, we investigated how LPD differs from random dropout. As indicated in Tab. 6, LPD increasingly drops points as the distance increases. Fog tends to obscure objects that are further away more significantly. Therefore, this demonstrates that LPD effectively represents point drop due to occlusions such as fog, which it is designed to simulate. Consequently, LPD is a more reasonable point dropout method to deal with adverse weather conditions than random point drop.
Table 7: LiDAR segmentation results (mIoU) with the various architecture and datasets. All models trained with SemanticKITTI.
@@
Experiments on Various Input Data Representation, Models and Datasets. We evaluated our method using different data representations, architectures, and datasets, specifically: (1) projection-based CENet and (2) voxel-based SPVCNN for backbones, and (3) the synthetic corruption dataset SemanticKITTI-C [12]. Tab. 7 shows our results.
SemanticKITTI-C has a unique data distribution with more samples in classes, such as sidewalks and vegetation, than SemanticSTF. Our method consistently improves performance across various datasets and architectures. Specifically, for the SemanticSTF dataset, our method increases the mIoU of CENet by +7.8, SPVCNN by +10.3, and MinkUNet by +8.1 mIoU. Similarly, for the SemanticKITTI-C dataset, the improvements are +3.9 for CENet, +0.4 for SPVCNN, and +5.6 mIoU for MinkUNet. These consistent enhancements across different input data representations, models, and datasets highlight the robustness and generalizability of our approach.
Table 8: Tabel R.3. Comparison with CosMix and UniMix. DA and DG denote domain adaptation and domain generalization.
Methods mIoU
CoSMix(DA) 28.4
UniMix(DA) 37.2
UniMix(DG) 31.4
MinkUnet18+Ours 36.3
Comparison with Domain Generalization and Adaptation Methods. CosMix [26] and UniMix [36] offer unsupervised domain adaptation (DA) techniques, with UniMix also providing domain generalization (DG) techniques. Domain generalization methods train models using only source domain data, while unsupervised domain adaptation methods train models with unlabeled target domain data. Our method, however, exclusively uses source data, making it fundamentally more challenging than UDA and not directly comparable. As shown in 8, our method outperforms CosMix and UniMix (DG) by +7.9 and +4.9 mIoU, respectively. This result demonstrates that our proposed method effectively represents adverse weather conditions without real weather-conditioned data. Additionally, our method performs comparably to UniMix (DA) despite not using target domain data as DA does. Because UDA can complement data augmentations, combining our method with UniMix could further enhance performance.
Table 9: Hyperparameter ablation in SemanticKITTI-to-SemanticSTF. The values we chose are highlighted in bold.
Selective Jittering
DSJ range mIoU ASJ $\Delta\theta$ mIoU Gaussian $\sigma$ mIoU
$\left[2,5\right]$ 39.3 $\frac{1}{2}\pi$ 38.1 0.001 37.5
$\left[3,6\right]$ 38.8 $\frac{3}{4}\pi$ 39.9 0.005 38.4
$\left[4,7\right]$ 37.8 $\pi$ 39.5 0.01 39.5
[5,8] 39.5 $\frac{5}{4}\pi$ 40.0 0.05 37.3
$\left[6,9\right]$ 37.8 $\frac{3}{2}\pi$ 39.3 0.1 37.9
deviation 0.8 deviation 0.2 deviation 1.3
Learable Point Drop
Batch size mIoU Discount $\gamma$ mIoU Decay $\epsilon$ mIoU
8 37.1 0.5 37.8 100 38.5
16 37.1 0.8 38.4 500 38.7
32 39.5 0.9 38.4 1000 39.5
64 38.2 0.99 39.5 2000 37.3
128 37.0 0.999 38.8 3000 40.3
deviation 1.7 deviation 0.9 deviation 0.7
Augmentation Configurations and Parameter Ablation. (1) DSJ adds noise to points along discretized depth levels. The number of depth levels, randomly selected within a specific range, is a hyperparameter. More depth levels improved rain resistance. (2) ASJ perturbs points along angles, with the maximum angular range $\footnotesize{\Delta\theta}$ as a hyperparameter. A larger angle range improved dense fog resistance. (3) Gaussian noise determines the jittering severity in SJ with its standard deviation $\sigma$ as a hyperparameter. Higher $\sigma$ improved resistance to rain and snow; lower $\sigma$ to fog. (4) LPD involves a batch size, discount rate $\gamma$, and exploration decay $\epsilon$ as hyperparameters related to DQN learning. Higher $\gamma$ improved fog resistance. Ablation studies in Tab. 9 show that we achieve generalized performance, less sensitive to hyperparameter choices.

### v2

@@ Methods
Figure 5: Qualitative results of LPD from validation set of SemanticKITTI. It can be observed that the point drop facilitated by the LPD occurs not just locally but extends across all points. While LPD effectively mimics the occlusions and sparse point distributions seen in real adverse weather, it does not replicate the effect of wet ground surfaces.
Experiments on Various Input Data Representation, Models and Datasets. We evaluated our method using different data representations, architectures, and datasets, specifically: (1) projection-based CENet and (2) voxel-based SPVCNN for backbones, and (3) the synthetic corruption dataset SemanticKITTI-C [11]. Tab.  4 shows our results.
SemanticKITTI-C has a unique data distribution with more samples in classes, such as sidewalks and vegetation, than SemanticSTF. Our method consistently improves performance across various datasets and architectures. Specifically, for the SemanticSTF dataset, our method increases the mIoU of CENet by +7.8, SPVCNN by +10.3, and MinkowskiNet by +8.1 mIoU. Similarly, for the SemanticKITTI-C dataset, the improvements are +3.9 for CENet, +0.4 for SPVCNN, and +5.6 mIoU for MinkowskiNet. These consistent enhancements across different input data representations, models, and datasets highlight the robustness and generalizability of our approach.
Table 4: LiDAR segmentation results (mIoU) with the various architecture and datasets. All models were trained with SemanticKITTI.
@@ 6 Conclusion
Acknowledgements. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-02068, Artificial Intelligence Innovation Hub and RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)), IITP grant funded by the Korea government(MSIT) and KEIT grant funded by the Korea government(MOTIE) (No.2022-0-00680, No.2022-0-01045), Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the MSIP (NRF-2022R1A2C3011154, RS-2023-00219019, RS-2023-00240135) and Samsung Electronics Co., Ltd (IO230508-06190-01).